OpenStudy (anonymous):

Hi How can I inverse these two equations: 1/ f(x)=2−e^2x 2/ f(x)=−8-(4/x) Thanks,

zepdrix (zepdrix):

\[Let \qquad f(x)=y\]To find the inverse function, we'll swap x and y in the equation.\[y=2-e^{2x} \rightarrow \qquad x=2-e^{2y}\] And now we want to solve for y :D Hmmm.

zepdrix (zepdrix):

\[\huge e^{2y}=2-x\]

zepdrix (zepdrix):

Now we need to get the y out of the exponent, hmmm do you know how to do that? :D I suppose we could take the natural log of both sides. \[\huge \ln ( e^{2y})=\ln ( 2-x)\]

zepdrix (zepdrix):

Next, we want to apply this rule of logarithms to get the y out of the log. \[\huge \log(a^b)=b \log (a)\]

zepdrix (zepdrix):

\[\huge (2y) \ln (e) = \ln (2-x)\]

zepdrix (zepdrix):

Andddd do you remember what the natural log of e is? It's basically asking, e to what power gives us e? e to the first power yes? :O so ln e = 1 If you're confused about that, maybe just try punching it into your calculator ^^ it's a good one to remember though.

zepdrix (zepdrix):

\[\huge 2y= \ln (2-x)\]

zepdrix (zepdrix):

\[\huge y=\frac{\ln (2-x)}{2}\] \[\huge f^{-1}(x)=\frac{\ln (2-x)}{2}\]

zepdrix (zepdrix):

Confused about anything in there? :O it's quite a few steps to get through! :D

OpenStudy (anonymous):

Thank you zepdrix :) that's really helpful. I was trying many many time and every time i get this: ln(x+2)/2

zepdrix (zepdrix):

Ah, yah i kinda flew through that step :3 hopefully you see what's going on. \[\large x=2-e^{2y}\]\[\large x-2=-e^{2y}\]Multiplying both sides by -1 gives:\[\large 2-x=e^{2y}\]

OpenStudy (anonymous):

yup :D i get it now. Can you help me with the 2nd equation plz?

zepdrix (zepdrix):

\[\large y=-8-\frac{4}{x}\]Taking the inverse gives us:\[\large x=-8-\frac{4}{y}\]

zepdrix (zepdrix):

\[\large x+8=-\frac{4}{y}\]Multiplying both sides by y gives us:\[\large y(x+8)=-4\]

zepdrix (zepdrix):

Dividing both sides by (x+8) gives us:\[\large y=\frac{-4}{x+8}\] \[\huge f^{-1}(x)=\frac{-4}{x+8}\]

OpenStudy (anonymous):

omg thank you. I was doing it wrong every time by multiplying y in x and then 8 and getting crazy answers! Sounds easy now, Thanks again :)

zepdrix (zepdrix):

hah :D